5x+3y+2z=-4
-5x-4y-2z=7
4x+2y+2z=-2
Solve with substitution or elimination

Answer by Alan3354(69443)

5x+3y+2z=-4
-5x-4y-2z=7
4x+2y+2z=-2
-------------
Add the middle eqn to the others to eliminate the z terms.
Then you have a system in x & y.
------------
I don't need the practice, but I'll check your work.